# A comprehensive study of the relations between the properties of planetary systems and the chemical compositions of their host stars

### v1

@@ II.4 Stellar Evolutionary Parameters
We determined the evolutionary parameters for our stars using the version 1.3 of PARAM (dasilva06)444https://stev.oapd.inaf.it/cgi-bin/param_1.3. The code employs a Bayesian statistical framework to estimate masses, radii and ages through a comparison of observational parameters with a grid of isochrones. We chose version 1.1 of the grid of PARSEC isochrones (bressan12) and kept the default options for the Bayesian priors. We also provided the following input parameters: , [Fe/H], parallax and V magnitude. The effective temperatures were taken from martinez19 and the metallicities from ghezzi21. The parallaxes were retrieved from Gaia Data Release 3 (DR3; gaia23), but two stars (KOI-4098 and KOI-4580) did not have available values. The V magnitudes were calculated from the G magnitudes and colors following Costa-Almeida et al. (in prep.). These magnitudes were then corrected for extinction using values calculated with the StarHorse code (queiroz18; anders19; anders22) and available in the contributions to Gaia Early Data Release 3 (EDR3; gaia21) for 549 stars and Gaia Data Release 2 (DR2; gaia18) for 9 stars555https://gaia.aip.de/cms/data/contrib/. For one star, was taken from the Galactic Dust Reddening and Extinction tool on the NASA/IPAC Infrared Science Archive (IRSA)666https://irsa.ipac.caltech.edu/applications/DUST/, adopting the value calculated with the extinction map of schlafly11. PARAM was unable to find a solution for eight stars. The masses, radii and ages for the remaining 551 stars are shown in Table 1.
@@ III.1 Chemical abundance distributions
The distributions of the chemical abundances determined in this work are shown as blue histograms in Figure 2, and their median values are shown as blue solid lines. We can see that our sample, which is composed only of planet hosts, has higher median abundances than the Sun for all elements. This is probably a consequence of the higher metallicity of our sample, having a median and MAD of 0.06 0.11 dex. As the distributions for all species but Cr I, Cr II and Mn are non-Gaussian according to a Shapiro-Wilk test (adopting the limit -value 0.001), we calculated the median values and they range from 0.05 (Mn) to 0.15 dex (V).
To check if these distributions are consistent with those from the solar neighborhood, we compared them with the ones obtained by luck17; luck18 from the analysis of a sample of 1777 stars in a “local region” within 100 pc of the Sun. Recall that stars in our sample, on the other hand, have distances ranging from 60 to 1800 pc. Since they used different solar abundances (scott15a; scott15b; grevesse15), we put their values on the same reference scale we adopted (asplund09) using the relation [X/H]rescaled = [X/H]L17+L18 + A(X) - A(X), where [X/H]L17+L18 are the abundances from luck17; luck18, A(X) are the absolute solar abundances from scott15a; scott15b and grevesse15 and A(X) are the absolute solar abundances from asplund09 for a given element X. As their sample spans wider intervals for the atmospheric parameters, we performed the following cuts in their sample: 4750 K 6350 K, 3.40 dex 4.90 dex and -0.60 dex [Fe/H] 0.50 dex. The more restricted sample contains 1368 stars, although not all abundances are available for all stars.
The abundance distributions from luck17; luck18 are represented by red histograms in Figure 2 and we can see a general good overlap between the two samples, but ours does not contain as many stars with lower abundances. Following ghezzi21, we investigated whether these samples are different using the Cucconi statistical test777We used the Python implementation developed by Grzegorz Mika and available at https://github.com/GrzegorzMika/NonParStat. We chose the option “bootstrap” for the method and 105 replications.. It simultaneously compares the central tendency and variance of two independent samples and is considered one of the most sensitive nonparametric tests to jointly evaluate differences in location and scale (marozzi13). As for Fe, the two sets of stars are statistically different ( 0.001) for all species, except Mg, Si, Sc II and Cr I. This result reveals that our sample is not representative of the solar neighborhood and this is probably a consequence of the higher metallicities of its stars.
@@ III.2.1 Possible correlations with stellar parameters
We investigated the behavior of the abundances as a function of and through weighted linear fits and found for and for . Despite these low correlation coefficients, we note that, according to Spearman and Pearson tests, we find significant correlations for the abundances of Na, Mg, Al, Ca, Ti I, V, Cr I, Cr II, Mn and Cu in the case of and for the abundances of Na, Ti I, V, Cr I, Cr II, Mn, Co and Cu in the case of .
Trends of increasing abundances as a function of decreasing have been observed in multiple previous works (e.g., adibekyan12a; luck17; luck18; brewer18). An intriguing case is Ti I, which in this study has a good number of lines with moderate to large EWs. We investigated whether this trend could be caused by blended lines, which we evaluated through spectral synthesis and the catalog of heiter21, but their removal did not improve the results. We also checked if non-LTE effects could be the cause of the trend, and although they may be significant for Ti lines, the non-LTE corrections for the parameter space of our sample are not high enough to explain the observed trend (e.g., mallinson22; mallinson24).
An additional test is to compare the abundances obtained from different ionization stages of the same element. For Ti, we confirm the increase in the Ti I abundances with decreasing and a similar behavior is observed for Sc I and Cr II. We note that Cr II and Sc I have only seven and four lines, respectively. Together with the typically low S/N values of our spectra, blendings could explain the increasing discrepancies towards lower effective temperatures. For Sc, we also see an increase in the differences for higher effective temperatures, which can be explained by the weakness of the Sc I lines for hotter stars coupled with the typical low S/N values of our spectra. Despite these trends, there is an overall good agreement between the two sets of abundances for each element. The median and MAD values of [X I/H] - [X II/H] are 0.013 0.079 dex for Sc, -0.020 0.050 dex for Ti and -0.045 0.053 dex for Cr. We note that there are no significant trends of the differences [X I/H] - [X II/H] with .
@@ III.2.2 Tests with the solar spectrum
The spectra analyzed in this work have a wide range of S/N values (10 – 300), and here we describe the tests we performed with the solar-proxy spectrum to check the reliability of our abundances. We first analyzed the spectrum of sunlight reflected off Vesta from ghezzi18, which was also observed with the HIRES spectrograph. We measured its S/N value for 143 apparent continuum regions by dividing the average normalized flux in each region by the standard deviation. We then performed seven rounds of 2-clippings until convergence was achieved and found a mean value of S/N = 307, which is close to the value of 315 at 5500 Å. Using the same methodology applied in martinez19 and ghezzi21, we obtained the following atmospheric parameters: = 5792 16 K, = 4.47 0.06, [Fe/H] = 0.02 0.01 dex and = 1.014 0.030 km s-1. These are in excellent agreement with the canonical solar values.
Adopting these parameters and following the methodology described in Section II.3, we determined the chemical abundances for 16 species (13 elements, including three - Sc, Ti and Cr - with two ionization stages) in the Sun and these can be found in Table 1 and the left panel of Figure 3. We note that all abundances agree with the solar values from asplund09 within 2, except for Na, Si and Cr II. This is a consequence of the small internal uncertainties that result from the analysis of a high-quality spectrum of the Sun using solar log gf values. The largest absolute difference relative to the reference values adopted from the work of asplund09 is 0.062 dex for Cr II, with an average = 0.015 0.026 dex, i.e., in excellent agreement with the solar abundances.
In order to test the performance of our methodology for lower quality spectra, we added noise to the solar-proxy spectrum to simulate the following S/N values: 250, 200, 150, 100, 90, 80, 70, 60, 50, 40, 30, 20 and 10. Utilizing the same atmospheric parameters and methodology as above, we obtained new abundances and calculated the absolute differences relative to those from asplund09. The results are shown in the right panel of Figure 3. We can see that we are able to recover almost all the original abundances within 0.10 dex down to S/N = 30. The only exception is Sc I, which has only a few weak lines in the Sun. Note also that the lower S/N spectra do not necessarily have the larger differences because distinct lines are removed by the clipping procedures, affecting the final average abundances. Below S/N = 30, the average difference for all elements also gets larger than 0.05 dex.
@@ III.3 Comparison with the literature
The chemical abundances of the studied elements, except for Sc, Co and Cu, were previously determined for the CKS stars by brewer18 using the Spectroscopy Made Easy Software (SME; pv17) which is based on the spectral synthesis method. They analyzed the same spectra as we did but with a different methodology. Their results are relative to the solar abundances of gs98, and thus we applied offsets to put their values on the same reference scale we adopted (asplund09). In Figure 4, we compare our abundances with theirs for 489 stars in common. The distributions of residuals are not Gaussian (-value 0.001) according to a Shapiro-Wilk test and the absolute median differences are 0.05 dex, except for Mg and Mn. The latter has the maximum offset of 0.108 dex. We performed linear fits to residuals and found 0.12 and slopes between -0.158 and 0.053. The Pearson and Spearman tests show that the trends in the residuals are statistically significant for Ca and Mn.
@@ IV.1 Abundance patterns for the stellar sample
The abundances of our stars are shown with the canonical diagrams [X/Fe] versus [Fe/H] in Figure 5. We can see that almost all elements follow their expected behaviors, with iron-peak elements (such as Cr, Co and Ni) having an almost constant abundance around [X/Fe] = 0 and the -elements (Mg, Si, Ca and Ti) exhibiting an increase towards lower metallicities. The only exception is Mn, which shows a slight increase with [Fe/H], also seen by wilson22.
One interesting feature in Figure 5 is the possible presence of -enhanced stars (adibekyan11; adibekyan12a) in our sample. To quantitatively assess this, in Figure 6, we show the [/Fe] abundance ratio (average of Mg, Si and Ti, as in adibekyan11) versus [Fe/H] for our sample along with the separation between the low- and high- sequences estimated from Figure 1 of adibekyan11. For [Fe/H] -0.6, [/Fe] = 0.19; for [Fe/H] 0.0, [/Fe] = 0.07; and for -0.6 [Fe/H] 0.0, [/Fe] = -0.20 [Fe/H] + 0.07.
According to this separation, our sample of Kepler planet hosts has 76 stars on the high-[/Fe] sequence and 434 stars on the low-[/Fe] sequence. The age distributions for these two samples are shown in Figure 7. The ages of the high-[/Fe] stars range from 0.92 to 11.50 Gyr, with a median value of 7.0 2.5 Gyr. Low-[/Fe] stars have ages between 0.24 and 11.50 Gyr, with a lower median value of 4.6 1.8 Gyr, which is similar to the age of the Sun (4.5 Gyr). The Cucconi test shows that these two samples of stars on the low- and high- sequences have statistically different distributions of ages (-value = 10-5). The median distance of the high- stars is 742 143 pc, while the low-[/Fe] stars have a smaller median value of 575 187 pc. As for the age, the distributions of distances are statistically different (-value = 4 ). Therefore, high- stars are typically older and more distant than their low- counterparts, as generally expected for a population of thick versus thin disk of the Galaxy.
Since the abundances vary over time due to Galactic Chemical Evolution (GCE), as can be seen in Figure 8, we also investigated whether there are any correlations between their values and stellar ages. We performed weighted least squares (WLS) fits (using the uncertainties on the abundances as the weights) and determined the coefficients and for the equation [X/Fe] = (aX Age) + bX for each element X (see Table 3), where [X/Fe] = [X/H] - [Fe/H]. According to Pearson and Spearman tests, we have statistically significant correlations for Mg, Al, Si, Ca, Sc II, Ti I, Ti II, Mn and Co. The stronger relation is observed for Mg, which shows a slope of 0.014 dex Gyr-1. In Figure 8, we can see that the relations determined in this work are in good agreement with those obtained by nissen17, bedell18 and liu20, despite the differences between the samples and methodologies adopted by the different studies.
| Species | a | b | -value | -value |
| (dex Gyr-1) | (dex) | (Spearman) | (Pearson) | |
@@ IV.2 Stellar abundances as a function of exoplanet properties
The 510 stars in our sample host 726 planets: 367 super-Earths (SE), 318 sub-Neptunes (SN), 23 sub-Saturns (SS) and 18 Jupiters (JP). In Figure 9, we show the derived stellar abundances [X/H] (shown as filled blue circles) as a function of the exoplanetary radii, with the latter taken from martinez19. We note that in this figure the abundances from hosts of multiplanetary systems are plotted individually and thus appear more than one time. We find that the median values (shown as filled red circles) are higher for stars with large planets (SS or JP) relative to those hosting small planets (SE or SN) for all elements and this result is consistent with the one found by ghezzi21 for Fe. The minimum and maximum differences between the median abundances are 0.055 dex for Ca and 0.155 dex for Cu, respectively. The median difference for all elements is 0.123 0.008 dex. The Pearson test shows that there are significant correlations (-value 0.001) for Na, Mg, Al, Si, Sc I, Sc II, Ti I, Ti II, Cr I, Mn, Co and Ni. The Spearman test, on the other hand, returns -values 0.001 for Mg, Sc II, Ti II, Mn and Co. We also performed Cucconi tests and found that the distributions of abundances for stars with large or small planets are significantly different (-value 0.001) for all 16 species, except Ca, V and Cu.
@@
If we use [X/Fe] abundances instead, the median values are similar for the four classes of planets (SE, SN, SS and JP), suggesting that there are no significant correlations with planet size. This is confirmed by Spearman and Pearson tests. These results agree with those from wilson22 for the elements analyzed in both works (Mg, Al, Si, Ca, Mn and Ni), except for [Mn/Fe], for which they find a correlation and we do not. However, wilson22 consider that their result for Mn was probably caused by a strong underlying correlation between [Mn/Fe] and [Fe/H] in the APOGEE abundance results (which we also observe in our sample). The absence of correlations between [X/Fe] and planet radii suggests that the trends observed for [X/H] reflect the underlying correlation between [Fe/H] and Rpl.
@@
We also investigated trends with orbital periods (taken from martinez19) by dividing our sample in stars with hot (P 10 days; N = 395) and warm (10 days P 100 days; N = 331) exoplanets. The median difference between stars with hot and warm planets is 0.050 0.009 for all elements, with a minimum value of 0.019 for Sc I and a maximum value of 0.076 for Al. We find statistically significant correlations (-value 0.001) for Ca and Cu in the Pearson test and for all species except Mg, Sc I and Ti II in the Spearman test. Moreover, the Cucconi tests reveal that the distributions of abundances for stars with hot or warm planets are significantly different (-value 0.001) for all 16 species, except Mg, Sc I, Sc II, Ti II and Ni. As before, if we use [X/Fe] abundances instead, we do not find -value 0.001 for any of the species. This result also agrees with the conclusions of wilson22 for all elements in common (Mg, Al, Si, Ca, Mn and Ni) and supports the hypothesis the trends observed for [X/H] simply reflect a dependence between orbital period and [Fe/H].

### v2

@@ II.4 Stellar Evolutionary Parameters
We determined the evolutionary parameters for our stars using the version 1.3 of PARAM (da Silva et al., 2006)444https://stev.oapd.inaf.it/cgi-bin/param_1.3. The code employs a Bayesian statistical framework to estimate masses, radii, and ages through a comparison of observational parameters with a grid of isochrones. We chose version 1.1 of the grid of PARSEC isochrones (Bressan et al., 2012) and kept the default options for the Bayesian priors. We also provided the following input parameters: , [Fe/H], parallax, and V magnitude. The effective temperatures were taken from Martinez et al. (2019), and the metallicities were from Ghezzi et al. (2021). The parallaxes were retrieved from Gaia Data Release 3 (DR3; Gaia Collaboration et al., 2023), but two stars (KOI-4098 and KOI-4580) did not have available values. The V magnitudes were calculated from the G magnitudes and colors following E. Costa-Almeida et al. (2026, in preparation). These magnitudes were then corrected for extinction using values calculated with the StarHorse code (Queiroz et al., 2018; Anders et al., 2019, 2022) and available in the contributions to Gaia Early Data Release 3 (EDR3; Gaia Collaboration et al., 2021) for 549 stars and Gaia Data Release 2 (DR2; Gaia Collaboration et al., 2018) for 9 stars555https://gaia.aip.de/cms/data/contrib/. For one star, was taken from the Galactic Dust Reddening and Extinction tool on the NASA/IPAC Infrared Science Archive (IRSA)666https://irsa.ipac.caltech.edu/applications/DUST/, adopting the value calculated with the extinction map of Schlafly and Finkbeiner (2011). PARAM was unable to find a solution for eight stars. The masses, radii, and ages for the remaining 551 stars are shown in Table 1.
@@ III.1 Chemical abundance distributions
The distributions of the chemical abundances determined in this work are shown as blue histograms in Figure 2, and their median values are shown as blue solid lines. We can see that our sample, which is composed only of planet hosts, has higher median abundances than the Sun for all elements. This is probably a consequence of the higher metallicity of our sample, having a median and MAD of 0.06 0.11 dex. As the distributions for all species but Cr I, Cr II, and Mn are non-Gaussian according to a Shapiro-Wilk test (adopting the limit -value 0.001), we calculated the median values and they range from 0.05 (Mn) to 0.15 dex (V).
To check if these distributions are consistent with those from the solar neighborhood, we compared them with the ones obtained by Luck (2017, 2018) from the analysis of a sample of 1777 stars in a “local region” within 100 pc of the Sun. Recall that stars in our sample, on the other hand, have distances ranging from 60 to 1800 pc. Since they used different solar abundances (Scott et al., 2015b, a; Grevesse et al., 2015), we put their values on the same reference scale we adopted (Asplund et al., 2009) using the relation [X/H]rescaled = [X/H]L17+L18 + A(X) - A(X), where [X/H]L17+L18 are the abundances from Luck (2017, 2018), A(X) are the absolute solar abundances from Scott et al. (2015b, a) and Grevesse et al. (2015), and A(X) are the absolute solar abundances from Asplund et al. (2009) for a given element X. As their sample spans wider intervals for the atmospheric parameters, we performed the following cuts in their sample: 4750 K 6350 K, 3.40 dex 4.90 dex and -0.60 dex [Fe/H] 0.50 dex. The more restricted sample contains 1368 stars, although not all abundances are available for all stars.
The abundance distributions from Luck (2017, 2018) are represented by red histograms in Figure 2, and we can see a general good overlap between the two samples, but ours does not contain as many stars with lower abundances. Following Ghezzi et al. (2021), we investigated whether these samples are different using the Cucconi statistical test777We used the Python implementation developed by Grzegorz Mika and available at https://github.com/GrzegorzMika/NonParStat. We chose the option “bootstrap” for the method and 105 replications.. It simultaneously compares the central tendency and variance of two independent samples and is considered one of the most sensitive nonparametric tests to jointly evaluate differences in location and scale (Marozzi, 2013). As for Fe, the two sets of stars are statistically different ( 0.001) for all species, except Mg, Si, Sc II, and Cr I. This result reveals that our sample is not representative of the solar neighborhood and this is probably a consequence of the higher metallicities of its stars.
@@ III.2.1 Possible correlations with stellar parameters
We investigated the behavior of the abundances as a function of and through weighted linear fits and found for and for . Despite these low correlation coefficients, we note that, according to Spearman and Pearson tests, we find significant correlations for the abundances of Na, Mg, Al, Ca, Ti I, V, Cr I, Cr II, Mn, and Cu in the case of and for the abundances of Na, Ti I, V, Cr I, Cr II, Mn, Co, and Cu in the case of .
Trends of increasing abundances as a function of decreasing have been observed in multiple previous works (e.g., Adibekyan et al., 2012b; Luck, 2017, 2018; Brewer and Fischer, 2018). An intriguing case is Ti I, which in this study has a good number of lines with moderate to large EWs. We investigated whether this trend could be caused by blended lines, which we evaluated through spectral synthesis and the catalog of Heiter et al. (2021), but their removal did not improve the results. We also checked if non-LTE effects could be the cause of the trend, and although they may be significant for Ti lines, the non-LTE corrections for the parameter space of our sample are not high enough to explain the observed trend (e.g., Mallinson et al., 2022, 2024).
An additional test is to compare the abundances obtained from different ionization stages of the same element. For Ti, we confirm the increase in the Ti I abundances with decreasing , and a similar behavior is observed for Sc I and Cr II. We note that Cr II and Sc I have only seven and four lines, respectively. Together with the typically low S/N values of our spectra, blendings could explain the increasing discrepancies towards lower effective temperatures. For Sc, we also see an increase in the differences for higher effective temperatures, which can be explained by the weakness of the Sc I lines for hotter stars coupled with the typical low S/N values of our spectra. Despite these trends, there is an overall good agreement between the two sets of abundances for each element. The median and MAD values of [X I/H] - [X II/H] are 0.013 0.079 dex for Sc, -0.020 0.050 dex for Ti, and -0.045 0.053 dex for Cr. We note that there are no significant trends of the differences [X I/H] - [X II/H] with .
@@ III.2.2 Tests with the solar spectrum
The spectra analyzed in this work have a wide range of S/N values (10 – 300), and here, we describe the tests we performed with the solar-proxy spectrum to check the reliability of our abundances. We first analyzed the spectrum of sunlight reflected off Vesta from Ghezzi et al. (2018), which was also observed with the HIRES spectrograph. We measured its S/N value for 143 apparent continuum regions by dividing the average normalized flux in each region by the standard deviation. We then performed seven rounds of 2 clippings until convergence was achieved and found a mean value of S/N = 307, which is close to the value of 315 at 5500 Å. Using the same methodology applied in Martinez et al. (2019) and Ghezzi et al. (2021), we obtained the following atmospheric parameters: = 5792 16 K, = 4.47 0.06, [Fe/H] = 0.02 0.01 dex, and = 1.014 0.030 km s-1. These are in excellent agreement with the canonical solar values.
Adopting these parameters and following the methodology described in Section II.3, we determined the chemical abundances for 16 species (13 elements, including three - Sc, Ti and Cr - with two ionization stages) in the Sun and these can be found in Table 1 and the left panel of Figure 3. We note that all abundances agree with the solar values from Asplund et al. (2009) within 2, except for Na, Si and Cr II. This is a consequence of the small internal uncertainties that result from the analysis of a high-quality spectrum of the Sun using solar log gf values. The largest absolute difference relative to the reference values adopted from the work of Asplund et al. (2009) is 0.062 dex for Cr II, with an average = 0.015 0.026 dex, i.e., in excellent agreement with the solar abundances.
In order to test the performance of our methodology for lower-quality spectra, we added noise to the solar-proxy spectrum to simulate the following S/N values: 250, 200, 150, 100, 90, 80, 70, 60, 50, 40, 30, 20, and 10. Utilizing the same atmospheric parameters and methodology as above, we obtained new abundances and calculated the absolute differences relative to those from Asplund et al. (2009). The results are shown in the right panel of Figure 3. We can see that we are able to recover almost all the original abundances within 0.10 dex down to S/N = 30. The only exception is Sc I, which has only a few weak lines in the Sun. Note also that the lower S/N spectra do not necessarily have the larger differences because distinct lines are removed by the clipping procedures, affecting the final average abundances. Below S/N = 30, the average difference for all elements also gets larger than 0.05 dex.
@@ III.3 Comparison with the literature
The chemical abundances of the studied elements, except for Sc, Co, and Cu, were previously determined for the CKS stars by Brewer and Fischer (2018) using the Spectroscopy Made Easy Software (SME; Piskunov and Valenti, 2017) which is based on the spectral synthesis method. They analyzed the same spectra as we did but with a different methodology. Their results are relative to the solar abundances of Grevesse and Sauval (1998), and thus we applied offsets to put their values on the same reference scale we adopted (Asplund et al., 2009). In Figure 4, we compare our abundances with theirs for 489 stars in common. The distributions of residuals are not Gaussian (-value 0.001) according to a Shapiro-Wilk test, and the absolute median differences are 0.05 dex, except for Mg and Mn. The latter has the maximum offset of 0.108 dex. We performed linear fits to residuals and found 0.12 and slopes between -0.158 and 0.053. The Pearson and Spearman tests show that the trends in the residuals are statistically significant for Ca and Mn.
@@ IV.1 Abundance patterns for the stellar sample
The abundances of our stars are shown with the canonical diagrams [X/Fe] versus [Fe/H] in Figure 5. We can see that almost all elements follow their expected behaviors, with iron-peak elements (such as Cr, Co, and Ni) having an almost constant abundance around [X/Fe] = 0 and the -elements (Mg, Si, Ca, and Ti) exhibiting an increase toward lower metallicities. The only exception is Mn, which shows a slight increase with [Fe/H], also seen by Wilson et al. (2022).
One interesting feature in Figure 5 is the possible presence of -enhanced stars (Adibekyan et al., 2011, 2012b) in our sample. To quantitatively assess this, in Figure 6, we show the [/Fe] abundance ratio (average of Mg, Si, and Ti, as in Adibekyan et al. 2011) versus [Fe/H] for our sample along with the separation between the low- and high- sequences estimated from Figure 1 of Adibekyan et al. (2011). For [Fe/H] -0.6, [/Fe] = 0.19; for [Fe/H] 0.0, [/Fe] = 0.07; and for -0.6 [Fe/H] 0.0, [/Fe] = -0.20 [Fe/H] + 0.07.
According to this separation, our sample of Kepler planet hosts has 76 stars on the high-[/Fe] sequence and 434 stars on the low-[/Fe] sequence. The age distributions for these two samples are shown in Figure 7. The ages of the high-[/Fe] stars range from 0.92 to 11.50 Gyr, with a median value of 7.0 2.5 Gyr. Low-[/Fe] stars have ages between 0.24 and 11.50 Gyr, with a lower median value of 4.6 1.8 Gyr, which is similar to the age of the Sun (4.5 Gyr). The Cucconi test shows that these two samples of stars on the low- and high- sequences have statistically different distributions of ages (-value = 10-5). The median distance of the high- stars is 742 143 pc, while the low-[/Fe] stars have a smaller median value of 575 187 pc. As for the age, the distributions of distances are statistically different (-value = 4 ). Therefore, high- stars are typically older and more distant than their low- counterparts, as generally expected for a population of thick versus thin disks of the Galaxy.
Since the abundances vary over time due to Galactic Chemical Evolution (GCE), as can be seen in Figure 8, we also investigated whether there are any correlations between their values and stellar ages. We performed weighted least-squares (WLS) fits (using the uncertainties on the abundances as the weights) and determined the coefficients and for the equation [X/Fe] = (aX Age) + bX for each element X (see Table 3), where [X/Fe] = [X/H] - [Fe/H]. According to Pearson and Spearman tests, we have statistically significant correlations for Mg, Al, Si, Ca, Sc II, Ti I, Ti II, Mn, and Co. The stronger relation is observed for Mg, which shows a slope of 0.014 dex Gyr-1. In Figure 8, we can see that the relations determined in this work are in good agreement with those obtained by Nissen et al. (2017), Bedell et al. (2018), and Liu et al. (2020), despite the differences between the samples and methodologies adopted by the different studies.
| Species | a | b | -value (Spearman) | -value (Pearson) |
| (dex Gyr-1) | (dex) | |||
@@ IV.2 Stellar abundances as a function of exoplanet properties
The 510 stars in our sample host 726 planets: 367 super-Earths (SE), 318 sub-Neptunes (SN), 23 sub-Saturns (SS) and 18 Jupiters (JP). In Figure 9, we show the derived stellar abundances [X/H] (shown as filled blue circles) as a function of the exoplanetary radii, with the latter taken from Martinez et al. (2019). We note that in this figure the abundances from hosts of multiplanetary systems are plotted individually and thus appear more than one time. We find that the median values (shown as filled red circles) are higher for stars with large planets (SS or JP) relative to those hosting small planets (SE or SN) for all elements, and this result is consistent with the one found by Ghezzi et al. (2021) for Fe. The minimum and maximum differences between the median abundances are 0.055 dex for Ca and 0.155 dex for Cu, respectively. The median difference for all elements is 0.123 0.008 dex. The Pearson test shows that there are significant correlations (-value 0.001) for Na, Mg, Al, Si, Sc I, Sc II, Ti I, Ti II, Cr I, Mn, Co, and Ni. The Spearman test, on the other hand, returns -values 0.001 for Mg, Sc II, Ti II, Mn, and Co. We also performed Cucconi tests and found that the distributions of abundances for stars with large or small planets are significantly different (-value 0.001) for all 16 species, except Ca, V, and Cu.
@@
If we use [X/Fe] abundances instead, the median values are similar for the four classes of planets (SE, SN, SS, and JP), suggesting that there are no significant correlations with planet size. This is confirmed by Spearman and Pearson tests. These results agree with those from Wilson et al. (2022) for the elements analyzed in both works (Mg, Al, Si, Ca, Mn, and Ni), except for [Mn/Fe], for which they find a correlation, and we do not. However, Wilson et al. (2022) consider that their result for Mn was probably caused by a strong underlying correlation between [Mn/Fe] and [Fe/H] in the APOGEE abundance results (which we also observe in our sample). The absence of correlations between [X/Fe] and planet radii suggests that the trends observed for [X/H] reflect the underlying correlation between [Fe/H] and Rpl.
@@
We also investigated trends with orbital periods (taken from Martinez et al. 2019) by dividing our sample in stars with hot (P 10 days; N = 395) and warm (10 days P 100 days; N = 331) exoplanets. The median difference between stars with hot and warm planets is 0.050 0.009 for all elements, with a minimum value of 0.019 for Sc I and a maximum value of 0.076 for Al. We find statistically significant correlations (-value 0.001) for Ca and Cu in the Pearson test and for all species except Mg, Sc I, and Ti II in the Spearman test. Moreover, the Cucconi tests reveal that the distributions of abundances for stars with hot or warm planets are significantly different (-value 0.001) for all 16 species, except Mg, Sc I, Sc II, Ti II, and Ni. As before, if we use [X/Fe] abundances instead, we do not find -value 0.001 for any of the species. This result also agrees with the conclusions of Wilson et al. (2022) for all elements in common (Mg, Al, Si, Ca, Mn, and Ni) and supports the hypothesis the trends observed for [X/H] simply reflect a dependence between orbital period and [Fe/H].